OpenStudy (rosemarybaldwin7):

Solve the following system of equations. Express your answer as an ordered pair in the format (a,b), with no spaces between the numbers or symbols. 2x+5y=16 -5x-2y=2

OpenStudy (qwertty123):

Do you know how to solve these?

OpenStudy (rosemarybaldwin7):

its been a while

OpenStudy (aveline):

How about you start by multiplying the top equation by 5 and the bottom one by 2? That way you can cancel out the x's

OpenStudy (qwertty123):

Simplifying 2x + 5y = 16 Solving 2x + 5y = 16 Move all terms containing y to the left, all other terms to the right.

OpenStudy (qwertty123):

Add '-2x' to each side of the equation. 2x + -2x + 5y = 16 + -2x Combine like terms: 2x + -2x = 0 0 + 5y = 16 + -2x 5y = 16 + -2x Divide each side by '5'.

OpenStudy (qwertty123):

After you Divide by 5 you get?

OpenStudy (rosemarybaldwin7):

each side meaning which numbers? im sorry im just a bit confused.

OpenStudy (aveline):



OpenStudy (aveline):

Can you find out what y is?

OpenStudy (rosemarybaldwin7):

okay i understand now for that part , do i put the two equations together?

OpenStudy (aveline):

Doing things the hard way :/ 21y=84 y=4

OpenStudy (rosemarybaldwin7):

oops duh i shouldve known to do 84 divided by 21. okay so i have the y now um where do i start to solve for x?

OpenStudy (rosemarybaldwin7):

nvm i think x is 0

OpenStudy (rosemarybaldwin7):

nvm again it was -2. which i dont know how to get that

OpenStudy (aveline):

2x+5y=16 We know that y=4, so: 2x+20=16 Subtract 20 from each side 2x=-4 Divide by 2 x=-2

OpenStudy (rosemarybaldwin7):

duh omg thank you i didnt get it before.

OpenStudy (aveline):

Np